The remainder when 3^12 + 5^12 is divided by 13 is 
 
pls belp

Grade:12th pass

1 Answers

Arun
4 years ago
3^12=27^4=(13*2+1)^4 thus remainder when dividing by 13 is 1.
5^12=25^6=(13*2-1)^6. thus remainder when divided by 13 is 1.
Thus total remainder is 2.